# Generating multipartite nonlocality to benchmark quantum computers

### v2

@@ Abstract
We show that quantum computers can be used for producing large $n$-partite nonlocality, thereby providing a method to benchmark them. The main challenges to overcome are: (i) The interaction topology might not allow arbitrary two-qubit gates. (ii) Noise limits the Bell violation. (iii) The number of combinations of local measurements grows exponentially with $n$. To overcome (i), we point out that graph states that are compatible with the two-qubit connectivity of the computer can be efficiently prepared. To mitigate (ii), we note that, for specific graph states, there are $n$-partite Bell inequalities whose resistance to white noise increases exponentially with $n$. To address (iii) for any $n$ and any connectivity, we introduce an estimator that relies on random sampling. As a result, we propose a method for producing $n$-partite Bell nonlocality with unprecedented large $n$. This allows in return to benchmark nonclassical correlations regardless of the number of qubits or the connectivity. We test our approach by using a simulation for a noisy IBM quantum computer, which predicts $n$-partite Bell nonlocality for at least $n=24$ qubits.
@@ I Introduction
The term “$n$-partite nonlocality” refers to correlations between $n$ parties that cannot be explained by any local realistic model [1, 2]. It can be detected by a violation of a multipartite Bell inequality, which shows that at least one of the assumptions of a local realistic model is false [3, 4]. The experimental test of $n$-partite nonlocality is however challenging. The first and main reason is that it is very difficult to have physical systems with $n$ parts that can be prepared in a genuinely $n$-partite entangled quantum state [5] and on which specific local measurements can be performed on each of the $n$ individual parts. From this perspective, quantum computers offer a unique chance to realistically go to a large $n$ and test quantum theory. Quantum computers have dozens, hundreds, or even thousands of qubits which can, in principle, be prepared in arbitrary quantum states and then measured individually.
@@
The second motivation is to use the observed Bell violation as a benchmark to compare different quantum computers. Since, as far as we know, $n$-partite nonlocality is a phenomenon specific to quantum theory, one can think of using it to quantify the “quantumness” of the device that has produced it. This can be achieved by the fraction $D_{n}=Q_{n}/C_{n}$ of the maximal quantum value $Q_{n}$ and the classical bound $C_{n}$ [8, 9, 10, 5]. For our use-case, we stress that $D_{n}$ is related to the resistance of the violation to depolarisation noise. Moreover, $D_{n}$ is associated to the detection efficiency that is required to classically simulate the quantum nonlocality [8].
@@
A variety of benchmarks have been proposed to test the quality of quantum computers, i.e., the quantum volume or the cross-entropy benchmark [11, 12, 13]. Still, no universally accepted standard has been established. Current approaches do not fulfill the ideal requirements to be independent of the noise model and the hardware, to be not tied to one algorithm but still being predictive and scalable in practice [12]. The phenomenon of $n$-partite nonlocality is promising in this regard, as a Bell violation has an interpretation independent of the hardware and the specific noise. Observing a violation of a Bell inequality requires specific quantum states and measurements. Consequently, Bell inequalities can be used to certify both measurements and states [14], which makes them attractive for benchmarking. In contrast, entanglement witnesses certify a quantum state given some well-characterized measurements, which requires additional assumptions. In addition, a Bell test can be carried out in principle for all pure entangled states [15]. In this sense, using Bell inequalities for benchmarking does not rely on a specific algorithm to prepare a certain quantum state. The fraction $D_{n}$ grows with system size and we show that this facilitates the scaling of the Bell test to many qubits. Finally, an observed Bell violation can be used to lower bound the fidelity, which allows to predict the quality of other computations.
@@
Experimentally, violations of $n$-partite Bell inequalities have been observed in a variety of physical systems that are promising for quantum computing [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. So far, however, mostly systems with a relatively small number of parties $n$ have been considered. In an ion trap, violations of up to $n=14$ have been verified [16], whereas nonlocality has also been shown for $n=6$ with photons [17]. $n=3$ nonlocality has been verified on a NMR quantum simulator [26]. Finally, nonlocality has also been investigated in atomic ensembles [27] and in optical lattices [18]. The largest number of parties has been achieved with superconducting qubits with up to $n=57$ [24]. To limit the experimental resources, however, this reference considered Bell inequalities that do not show an exponential violation in $n$.
@@
The question remains why exponentially increasing nonlocality has not been verified with quantum computers before? There are, fundamentally, three challenges:
@@
(i) Typically, quantum computers can apply two-qubit gates only on some specific pairs of qubits. Hereafter, we will refer to the map that specifies these pairs as the connectivity of the quantum computer. A consequence of this limitation is that not all connectivities allow us to equally well prepare a $n$-qubit Greenberger-Horne-Zeilinger (GHZ) [28] state, which was the default option in [19, 16].
@@
(iii) The number of different combinations of local measurements (contexts) needed to test a Bell inequality increases with $n$. For example, in the case that there are two measurements per qubit, the number of contexts scales exponentially in $n$ and so does typically the number of terms needed to test the Bell inequality. Therefore, measuring all contexts becomes infeasible for large $n$.
In this work we show how to overcome or alleviate each of these three challenges. First, we will show that there is a natural solution to problem (i). For a given connectivity, we can focus on the graph states that are compatible with the connectivity graph. Graph states are a specific set of pure entangled states [29] and can be prepared by applying controlled-Z (CZ) gates on adjacent qubits in the graph. Thus, if we assume that CZ gates can be performed between connected qubits, graph states that correspond to subgraphs of the two-qubit connectivity can be readily prepared.
For $n$-qubit graph states, there are some general methods to obtain $n$-partite Bell inequalities [10, 30, 8, 31, 32]. Though, it is in general a hard task to find the optimal one (in the sense of resistance to noise of the violation), the optimal $n$-partite Bell inequalities in terms of the stabilizers have been identified for some graph states [8]. In particular, the optimal $n$-partite Bell inequalities associated to the GHZ and linear cluster (LC) state are known [8, 30]. These states correspond to the extreme cases of connectivity: On the one side, the GHZ state is easy to prepare when all qubits can be coupled to each other. On the other side, we consider the LC state that can be conveniently prepared on a quantum computer with minimal connectivity, i.e., the connectivity graph is a line. Cluster states have in addition the advantage to be more resistant to decoherence [33]. Quantum theory predicts that for the GHZ and LC state the ratio $D_{n}$ can be made arbitrarily large by increasing the number of qubits [5, 34, 30]. This means that, in theory, the resistance to noise of multipartite nonlocality grows exponentially with the number of particles. This helps to overcome (ii).
@@
We note however that quantum computers usually are not suited for a loophole-free Bell test. For example, ions are typically in the same trap and superconducting qubits on the same chip. It is thus not possible to rule out communication. However, in principle, the interactions can be tuned to minimize the cross-talk between the qubits. Hereafter, we will refer to this assumption as the no-cross-talk assumption and we will make it on the belief that quantum computers are the only way to investigate $n$-partite nonlocality with large $n$.
To explain the proposed method, we will introduce graph states that can be readily prepared on a given connectivity. Accordingly, we discuss examples of Bell inequalities associated to graph states that show an exponential scaling of the fraction $D_{n}$ in Sec. II. Afterwards we will explain in Sec. III the method to measure the Bell inequalities. As we propose to evaluate the Bell inequalities by random sampling in Sec. III.2, we first formulate the Bell test as a hypothesis test in Sec. III.1. After we discuss the sample complexity in Sec. III.3, we will apply the method exemplarily to the Bell inequalities of the GHZ and LC state in Sec. IV. These Bell inequalities cover the extreme cases of connectivities in quantum computers. In doing so, we consider actual architectures of current quantum computers. Finally, we include a simulation for the IBM Eagle quantum processor in Sec. V.
@@ II Bell inequalities with an exponential nonlocality
We start this section by discussing graph states, which are an important class of quantum states in quantum information theory [29, 38]. For our aim, they are specifically useful as the graph states that are compatible with the connectivity graph of a $n$-qubit quantum computer can be readily prepared. Suppose $G$ is a graph of $n$ vertices. For each vertex $i$ in $G$, we define a stabilizing operator $g_{i}$ by
@@
where $\mathcal{N}(i)$ is the neighbourhood of vertex $i$, i.e., all vertices that are connected to $i$. In the above definition, $X_{i}$, $Y_{i}$ and $Z_{i}$ denote the Pauli matrices acting on qubit $i$. The graph state |$\ket{G}$ that is associated with $G$ is the common eigenstate of all stabilizing operators with eigenvalue $+1$, i.e.,
@@
Figure 1: Three exemplary graph states: (a) For three qubits, the GHZ state and the LC state coincide. The star graph in (b) corresponds to the GHZ state of six qubits, whereas the line graph in (c) is associated to the LC state of six qubits.
Different graphs may lead to graph states that are connected by a local unitary (LU) transformation, which does not change the entanglement properties. A special class of LU transformations are local Clifford operations [29], that can be described by a graphical rule called local complementation. A local complementation in vertex $i$ transforms a graph $G=(V,E)$ into a graph $G^{\prime}=(V,E^{\prime})$ by inverting the neighbourhood $\mathcal{N}(i)$ of vertex $i$. For two vertices $j,k\in\mathcal{N}(i)$, if $(j,k)\in E$ then $(j,k)\notin E^{\prime}$ and vice versa. The set of vertices $V$ is unchanged.
In addition, graph states have also another benefit. For all graph states there is an associated Bell inequality known that is maximally violated by the graph state [10]. Specifically, for tree graphs, i.e., connected graphs that do not contain a cycle, the violation increases exponentially in the number of qubits $n$ [10], i.e., the ratio $D_{n}^{\text{}}$ increases exponentially in $n$. This facilitates the observation of a Bell violation. Suppose in an experiment the noisy graph state $\rho=\alpha\outerproduct{G}{G}+(1-\alpha)\mathbbm{1}/2^{n}$ is prepared. The expectation value of the Bell operator is $\langle{\mathcal{B}_{n}^{\text{}}}\rangle_{\rho}=\alpha Q_{n}^{\text{}}$. A violation is thus observed for
@@ III Methods
The Bell operators in Sec. II rely on a number of contexts that scales exponentially in the number of qubits $n$, i.e., there are in principle an exponential number of terms to measure. The number of observables thus quickly becomes infeasible. In this section, we introduce a general method to evaluate a Bell inequality by sampling random terms. The terms of the Bell inequality are picked according to a uniform probability distribution. This approach stands in line with previous schemes that use randomization to reduce the number of measurements, e.g., direct fidelity estimation [35, 36] or few-copy entanglement detection [37]. Finally, there exist different methods to assess the statistical strength of Bell tests [39, 40, 41, 42]. We gauge the significance of a violation with the help of the $p$-value. For this purpose, we start by formulating a Bell test as a hypothesis test.
@@ III.1 Hypothesis test
The expectation value on the right hand side however cannot be inferred exactly in an experiment. Rather the expectation value has to be estimated from multiple experimental repetitions. For this purpose it is useful to consider an unbiased estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$, which we denote by a hat. An estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ is a function of the experimental data and it is unbiased in case it reproduces the actual value in expectation, i.e., $\mathbb{E}[\langle{\hat{\mathcal{B}^{\text{}}}}\rangle]=\langle{\mathcal{B}^{\text{}}}\rangle$.
Due to the finite statistics, however, the estimate $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ fluctuates. There is thus a non-zero probability to observe a violation of the Bell inequality although the actual state does not violate it. To quantify the probability that an observed violation is only due to statistical fluctuations, we formulate the Bell test as a hypothesis test. The hypotheses are
  • Null hypothesis $\mathbf{H_{0}}$: The measurement outcomes can be described by a local hidden variable (LHV) model,
@@
To gauge whether the observed data is in contradiction with the null hypothesis $H_{0}$, we look at the $p$-value. The $p$-value is defined as the probability to observe an estimate at least as large as some value $\theta$ in case $H_{0}$ is true, i.e.,
@@
But the $p$-value is hard to calculate since the probability depends on the probability distribution of the estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ which is unknown. We can, however, upper bound the $p$-value. A local theory can at most reach a value of $\langle{\mathcal{B}^{\text{}}}\rangle=\mathbb{E}[\langle{\hat{\mathcal{B}^{\text{}}}}\rangle]=C$. Thus, in case $H_{0}$ is true the estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ has to exceed its mean by at least $t=\theta-C$ if a violation ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\theta>C$ is observed. This is sketched in Fig. 2. We consequently obtain the upper bound
@@
In the following we use Hoeffding’s inequality [43] to upper bound the right hand side of Eq. (11). Hoeffding’s inequality is a large deviation bound that typically involves the number of repetitions. This allows us to connect the number of repetitions to the $p$-value. Finally, we say that an observed result has a confidence level of $\gamma=1-p$.
Figure 2: Upper bound of the $p$-value. To observe a value of $\theta>C$ even though the state obeys the classical bound, the estimator $\langle{\hat{\mathcal{B}}}\rangle$ has to exceed its mean by at least $t$.

### v3

@@ Abstract
We show that quantum computers can be used for producing large $n$-partite nonlocality, thereby providing a method to benchmark them. The main challenges to overcome are as follows: (i) The interaction topology might not allow arbitrary two-qubit gates. (ii) Noise limits the Bell violation. (iii) The number of combinations of local measurements grows exponentially with $n$. To overcome (i), we point out that graph states that are compatible with the two-qubit connectivity of the computer can be efficiently prepared. To mitigate (ii), we note that for specific graph states, there are $n$-partite Bell inequalities whose resistance to white noise increases exponentially with $n$. To address (iii) for any $n$ and any connectivity, we introduce an estimator that relies on random sampling. As a result, we propose a method for producing $n$-partite Bell nonlocality with unprecedented large $n$. This allows one, in return, to benchmark nonclassical correlations regardless of the number of qubits or the connectivity. We test our approach by using a simulation for a noisy IBM quantum computer, which predicts $n$-partite Bell nonlocality for at least $n=24$ qubits.
@@ I Introduction
The term “$n$-partite nonlocality” refers to correlations between $n$ parties that cannot be explained by any local realistic model [1, 2]. It can be detected by a violation of a multipartite Bell inequality, which shows that at least one of the assumptions of a local realistic model is false [3, 4]. The experimental test of $n$-partite nonlocality is, however, challenging. The first and main reason is that it is very difficult to have physical systems with $n$ parts that can be prepared in a genuinely $n$-partite entangled quantum state [5] and on which specific local measurements can be performed on each of the $n$ individual parts. From this perspective, quantum computers offer a unique chance to realistically go to a large $n$ and test quantum theory. Quantum computers have dozens, hundreds, or even thousands of qubits which can, in principle, be prepared in arbitrary quantum states and then measured individually.
@@
The second motivation is to use the observed Bell violation as a benchmark to compare different quantum computers. Since, as far as we know, $n$-partite nonlocality is a phenomenon specific to quantum theory, one can think of using it to quantify the “quantumness” of the device that has produced it. This can be achieved by the fraction $D_{n}=Q_{n}/C_{n}$ of the maximal quantum value $Q_{n}$ and the classical bound $C_{n}$ [8, 9, 10, 5]. For our use-case, we stress that $D_{n}$ is related to the resistance of the violation to depolarization noise. Moreover, $D_{n}$ is associated to the detection efficiency that is required to classically simulate the quantum nonlocality [8].
@@
A variety of benchmarks have been proposed to test the quality of quantum computers, i.e., the quantum volume or the cross-entropy benchmark [11, 12, 13]. Still, no universally accepted standard has been established. Current approaches do not fulfill the ideal requirements to be independent of the noise model and the hardware: to not be tied to one algorithm but still being predictive and scalable in practice [12]. The phenomenon of $n$-partite nonlocality is promising in this regard, as a Bell violation has an interpretation independent of the hardware and the specific noise. Observing a violation of a Bell inequality requires specific quantum states and measurements. Consequently, Bell inequalities can be used to certify both measurements and states [14], which makes them attractive for benchmarking. In contrast, entanglement witnesses certify a quantum state given some well-characterized measurements, which requires additional assumptions. In addition, a Bell test can be carried out, in principle, for all pure entangled states [15]. In this sense, using Bell inequalities for benchmarking does not rely on a specific algorithm to prepare a certain quantum state. The fraction $D_{n}$ grows with system size and we show that this facilitates the scaling of the Bell test to many qubits. Finally, an observed Bell violation can be used to lower bound the fidelity, which allows one to predict the quality of other computations.
@@
Experimentally, violations of $n$-partite Bell inequalities have been observed in a variety of physical systems that are promising for quantum computing [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. So far, however, mostly systems with a relatively small number of parties, $n$, have been considered. In an ion trap, violations of up to $n=14$ have been verified [16], whereas nonlocality has also been shown for $n=6$ with photons [17]. $n=3$ nonlocality has been verified on a nuclear magnetic resonance (NMR) quantum simulator [26]. Finally, nonlocality has also been investigated in atomic ensembles [27] and in optical lattices [18]. The largest number of parties has been achieved with superconducting qubits with up to $n=57$ [24]. To limit the experimental resources, however, this reference considered Bell inequalities that do not show an exponential violation in $n$.
@@
The question remains why exponentially increasing nonlocality has not been verified with quantum computers before. There are, fundamentally, three challenges:
@@
(i) Typically, quantum computers can apply two-qubit gates only on some specific pairs of qubits. Hereafter, we will refer to the map that specifies these pairs as the connectivity of the quantum computer. A consequence of this limitation is that not all connectivities allow us to equally well prepare an $n$-qubit Greenberger-Horne-Zeilinger (GHZ) [28] state, which was the default option in [19, 16].
@@
(iii) The number of different combinations of local measurements (contexts) needed to test a Bell inequality increases with $n$. For example, in the case that there are two measurements per qubit, the number of contexts scales exponentially in $n$ and, typically, so does the number of terms needed to test the Bell inequality. Therefore, measuring all contexts becomes infeasible for large $n$.
In this work, we show how to overcome or alleviate each of these three challenges. First, we will show that there is a natural solution to problem (i). For a given connectivity, we can focus on the graph states that are compatible with the connectivity graph. Graph states are a specific set of pure entangled states [29] and can be prepared by applying controlled-Z (CZ) gates on adjacent qubits in the graph. Thus, if we assume that CZ gates can be performed between connected qubits, graph states that correspond to subgraphs of the two-qubit connectivity can be readily prepared.
For $n$-qubit graph states, there are some general methods to obtain $n$-partite Bell inequalities [10, 30, 8, 31, 32]. However, it is, in general, a hard task to find the optimal one (in the sense of resistance to noise of the violation); the optimal $n$-partite Bell inequalities in terms of the stabilizers have been identified for some graph states [8]. In particular, the optimal $n$-partite Bell inequalities associated to the GHZ and linear cluster (LC) state are known [8, 30]. These states correspond to the extreme cases of connectivity: On the one side, the GHZ state is easy to prepare when all qubits can be coupled to each other. On the other side, we consider the LC state that can be conveniently prepared on a quantum computer with minimal connectivity, i.e., the connectivity graph is a line. Cluster states have, in addition, the advantage to be more resistant to decoherence [33]. Quantum theory predicts that for the GHZ and LC states, the ratio $D_{n}$ can be made arbitrarily large by increasing the number of qubits [5, 34, 30]. This means that in theory, the resistance to noise of multipartite nonlocality grows exponentially with the number of particles. This helps to overcome (ii).
@@
We note, however, that quantum computers usually are not suited for a loophole-free Bell test. For example, ions are typically in the same trap and superconducting qubits on the same chip. It is thus not possible to rule out communication. However, in principle, the interactions can be tuned to minimize the crosstalk between the qubits. Hereafter, we will refer to this assumption as the no-crosstalk assumption and we will make it on the belief that quantum computers are the only way to investigate $n$-partite nonlocality with large $n$.
To explain the proposed method, we will introduce graph states that can be readily prepared on a given connectivity. Accordingly, we discuss examples of Bell inequalities associated to graph states that show an exponential scaling of the fraction $D_{n}$ in Sec. II. Afterwards, we will explain in Sec. III the method to measure the Bell inequalities. As we propose to evaluate the Bell inequalities by random sampling in Sec. III.2, we first formulate the Bell test as a hypothesis test in Sec. III.1. After we discuss the sample complexity in Sec. III.3, we will apply the method exemplarily to the Bell inequalities of the GHZ and LC states in Sec. IV. These Bell inequalities cover the extreme cases of connectivities in quantum computers. In doing so, we consider actual architectures of current quantum computers. Finally, we include a simulation for the IBM Eagle quantum processor in Sec. V.
@@ II Bell inequalities with an exponential nonlocality
We start this section by discussing graph states, which is an important class of quantum states in quantum information theory [29, 38]. For our aim, they are specifically useful as the graph states that are compatible with the connectivity graph of an $n$-qubit quantum computer can be readily prepared. Suppose $G$ is a graph of $n$ vertices. For each vertex $i$ in $G$, we define a stabilizing operator $g_{i}$ by
@@
where $\mathcal{N}(i)$ is the neighborhood of vertex $i$, i.e., all vertices that are connected to $i$. In the above definition, $X_{i}$, $Y_{i}$ and $Z_{i}$ denote the Pauli matrices acting on qubit $i$. The graph state |$\ket{G}$ that is associated with $G$ is the common eigenstate of all stabilizing operators with eigenvalue $+1$, i.e.,
@@
Figure 1: Three exemplary graph states. (a) For three qubits, the GHZ state and the LC state coincide. The star graph in (b) corresponds to the GHZ state of six qubits, whereas the line graph in (c) is associated to the LC state of six qubits.
Different graphs may lead to graph states that are connected by a local unitary (LU) transformation, which does not change the entanglement properties. A special class of LU transformations is local Clifford operations [29], which can be described by a graphical rule called local complementation. A local complementation in vertex $i$ transforms a graph $G=(V,E)$ into a graph $G^{\prime}=(V,E^{\prime})$ by inverting the neighborhood $\mathcal{N}(i)$ of vertex $i$. For two vertices $j,k\in\mathcal{N}(i)$, if $(j,k)\in E$ then $(j,k)\notin E^{\prime}$, and vice versa. The set of vertices $V$ is unchanged.
In addition, graph states also have another benefit. For all graph states, there is a known associated Bell inequality that is maximally violated by the graph state [10]. Specifically, for tree graphs, i.e., connected graphs that do not contain a cycle, the violation increases exponentially in the number of qubits, $n$ [10], i.e., the ratio $D_{n}^{\text{}}$ increases exponentially in $n$. This facilitates the observation of a Bell violation. Suppose, in an experiment, that the noisy graph state $\rho=\alpha\outerproduct{G}{G}+(1-\alpha)\mathbbm{1}/2^{n}$ is prepared. The expectation value of the Bell operator is $\langle{\mathcal{B}_{n}^{\text{}}}\rangle_{\rho}=\alpha Q_{n}^{\text{}}$. A violation is thus observed for
@@ III Methods
The Bell operators in Sec. II rely on a number of contexts that scales exponentially in the number of qubits $n$, i.e., there are, in principle, an exponential number of terms to measure. The number of observables thus quickly becomes infeasible. In this section, we introduce a general method to evaluate a Bell inequality by sampling random terms. The terms of the Bell inequality are picked according to a uniform probability distribution. This approach stands in line with previous schemes that use randomization to reduce the number of measurements, e.g., direct fidelity estimation [35, 36] or few-copy entanglement detection [37]. Finally, there exist different methods to assess the statistical strength of Bell tests [39, 40, 41, 42]. We gauge the significance of a violation with the help of the $p$ value. For this purpose, we start by formulating a Bell test as a hypothesis test.
@@ III.1 Hypothesis test
The expectation value on the right-hand side, however, cannot be inferred exactly in an experiment. Rather, the expectation value has to be estimated from multiple experimental repetitions. For this purpose, it is useful to consider an unbiased estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$, which we denote by a hat. An estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ is a function of the experimental data and it is unbiased in case it reproduces the actual value in expectation, i.e., $\mathbb{E}[\langle{\hat{\mathcal{B}^{\text{}}}}\rangle]=\langle{\mathcal{B}^{\text{}}}\rangle$.
Due to the finite statistics, however, the estimate $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ fluctuates. There is thus a nonzero probability to observe a violation of the Bell inequality, although the actual state does not violate it. To quantify the probability that an observed violation is only due to statistical fluctuations, we formulate the Bell test as a hypothesis test. The hypotheses are as follows:
  • (a)
    Null hypothesis $\mathbf{H_{0}}$: The measurement outcomes can be described by a local hidden variable (LHV) model.
  • (b)
@@
To gauge whether the observed data is in contradiction with the null hypothesis $H_{0}$, we look at the $p$ value. The $p$ value is defined as the probability to observe an estimate at least as large as some value $\theta$ in case $H_{0}$ is true, i.e.,
@@
But the $p$ value is hard to calculate since the probability depends on the probability distribution of the estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$, which is unknown. We can, however, upper bound the $p$ value. A local theory can, at most, reach a value of $\langle{\mathcal{B}^{\text{}}}\rangle=\mathbb{E}[\langle{\hat{\mathcal{B}^{\text{}}}}\rangle]=C$. Thus, in case $H_{0}$ is true, the estimator $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ has to exceed its mean by at least $t=\theta-C$ if a violation ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\theta>C$ is observed. This is sketched in Fig. 2. We consequently obtain the upper bound
@@
In the following, we use Hoeffding’s inequality [43] to upper bound the right-hand side of Eq. (11). Hoeffding’s inequality is a large deviation bound that typically involves the number of repetitions. This allows us to connect the number of repetitions to the $p$ value. Finally, we say that an observed result has a confidence level of $\gamma=1-p$.
Figure 2: Upper bound of the $p$ value. To observe a value of $\theta>C$ even though the state obeys the classical bound, the estimator $\langle{\hat{\mathcal{B}}}\rangle$ has to exceed its mean by at least $t$.